OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 5y=x-7 4x-y=-10 A. (-8,3) B. (1/5,8) C. (-3,-2) D. (-1,-1)

OpenStudy (anonymous):

C is the answer...

OpenStudy (anonymous):

First get variables on same side. 5y - x = -7 and 4x-y = -10 then multiply top equation by 4 so x's will cancel in nxt step. Now you have 20y -4x = -28 +(4x - y) = -10 so 19y = -38 y = -2 now sub -2 back into equation to get 5(-2) = x-7 x = -3 So yes C is the answer, but in case you wanted to know how to solve the problem yourself next time I gave you the steps to do so if you need further explanation let me know.